OpenStudy (askme12345):

if joe has 6,000$ and wishes to invest it at 9% for 7 years how much will he have after 7 years, both are compounded continiously

OpenStudy (askme12345):

a) if kurt is competing with him and has the same amount of money but wishes to double it in 7 years, what rate should kurt have

OpenStudy (johnweldon1993):

Alright so we have a compounded continuoisly problem so \[\large A = Pe^{rt}\] P = 6000 r = .09 t = 7 \[\large A = 6000e^{.09 \times 7}\] Thats for problem 1 For problem 2...we have the same formula...except we dont have 'r' this time We know he wants double the amount he had to begin with....he wants 12000 dollars so A = 12000 P = 6000 t = 7 \[\large 12000 = 6000e^{7r}\] Can you solve these?

OpenStudy (askme12345):

so the answer for problem 1 would be 11265.66

OpenStudy (johnweldon1993):

Correct!

OpenStudy (askme12345):

not 100% sure how to solve part 2, i got up to 12000=6579798.95057^r

OpenStudy (askme12345):

do i do the square root

OpenStudy (johnweldon1993):

Not quite....for that problem we will use logarithms.... \[\large 12000 = 6000e^{7r}\] First divide both sides by 6000 what do you have?

OpenStudy (askme12345):

2000=e\[^{7r}\]

OpenStudy (askme12345):

sorry.... 2000=e^7r

OpenStudy (johnweldon1993):

hmmm 12000/6000? Looks like 2 to me :) right?

OpenStudy (askme12345):

ah sorry!! you are correct

OpenStudy (johnweldon1993):

No problem...so we have \[\large 2 = e^{7r}\] right?

OpenStudy (askme12345):

yup

OpenStudy (johnweldon1993):

Alright...now, what relationship is there between \(\large \ln\) and \(\large e\) ?

OpenStudy (askme12345):

they are inverse fuctions

OpenStudy (johnweldon1993):

Correct... Meaning...they cancel each other out...so if you took the 'ln' of both sides of this function...you would have \[\large ln(2) = ln(e^{7r})\] Now remember a rule \[\large lna^x = x \times ln(a)\] So now we have \[\large ln(2) = 7r \times ln(e)\] We know since ln and e are inverse functions...so \(\large ln(e)\) just = \(\large 1\) So we have \[\large ln(2) = 7r\] How would you solve for 'r' ?

OpenStudy (askme12345):

would you divide ln(2) by 7

OpenStudy (johnweldon1993):

^Indeed^ Did everything make sense though?

OpenStudy (askme12345):

Yes thank you SOOO MUCH!

OpenStudy (johnweldon1993):

No problem :)